Nora and Addison went shopping for new summer clothes. Nora bought 4 pairs of shorts and 7 tank tops for $102.50. Addison bought 2 pairs of shorts and 6 tank tops for $77.00. Find the costs of the shorts and tank tops.
Answer by 1998Baby(63)

4s+7t=102.5
2s+6t=77.0+> multiply bottem row by -2
______
you should get
4s+7t=102.5
-4s-12t=-154.0
than add down.
-5t=-51.5
t=$10.30 each
put t in any equation.
4s-12(10.3)=-154.0
4s-123.6=-154.0
4s=30.9
s=$7.725
4(7.725)+7(10.3)=102.5
30.9+72.1=$103 for Nora
2(7.725)+6(10.3)=77.0
15.45+61.8=77.25 for Addison
_______
sorry this is the best i could do